# When Bias Helps Learning: Bridging Initial Prejudice and Trainability

### v3

@@ 3 Connecting classification bias to the ordered phase
We report the proof of this theorem in App. E. The key result of Thm. 3.1 establishes a correspondence between the MF formulation and IGB, enabling signal propagation in wide networks to be described interchangeably using either framework. Intriguingly, this result makes no further assumption than the MF regime, and it remains valid for every architecture defined by Eq. (1). In MF theory, $q_{aa}^{(l)}$ and $q_{ab}^{(l)}$ are generally functions of the dataset $\mathcal{D}$, and therefore become random variables upon imposing a distribution over $\mathcal{D}$, as done in the IGB approach. Remarkably, these two quantities concentrate around their mean in the infinite-width and then -data limit, allowing their treatment as deterministic variables (equivalently through $\sigma^{2}_{y^{(l)}}$ and $\sigma^{2}_{\mu^{(l)}}$). On the one hand, this formal connection between MF and IGB frameworks enables a unified view, where predictive behaviour at initialization and trainability conditions are jointly entangled, enriching the classical MF picture, as will be discussed in Secs. 4 and 5. On the other, it extends the IGB framework from [pmlr-v235-francazi24a] to settings not previously analyzed (e.g., identifying prejudice and neutrality phases for Tanh activations directly from the MF phase diagram [schoenholz2016deep]).
In Fig. 2, we test the validity of Thm. 3.1 by plotting the correlation coefficient in function of the depth for ReLU and Tanh MLPs with $\sigma^{2}_{b}=0.1$ and $\sigma^{2}_{w}$ uniformly varying from the ordered to the chaotic phase analyzed in MF theory. We observe a good agreement between the curves obtained with the IGB approach (solid lines - computed via Eq. (6)) and the 90 $\%$ central confidence interval computed using the MF approach (shaded areas). The distribution of MF is very narrow around the IGB-computed values, corroborating the treatment of the signal variance and covariance as deterministic variables. As network’s width increases, the MF distributions of $q_{aa}^{(l)}$ and $q_{ab}^{(l)}$ become progressively more concentrated (see App. E). For ReLU, we observe that the correlation coefficient always converges to one, but the convergence rate is exponential in the ordered phase ($\sigma^{2}_{w}<2$) and follows a power-law in the chaotic phase ($\sigma^{2}_{w}>2$). For Tanh, the correlation coefficient converges to one in the ordered phase and to a lower value in the chaotic phase; in this case, we always observe an exponential convergence behaviour.
Figure 2: Convergence behaviour the correlation coefficient of ReLU and Tanh for a single MLP with width equal to 10 000 and depth 100. $\sigma^{2}_{b}=0.1$ and $\sigma^{2}_{w}$ varies uniformly from the ordered phase (blue) to the chaotic phase (red). The transition point is $\sigma^{2}_{w}=2.0$ for ReLU and close to it for Tanh. Scatter points indicate the asymptotic values. The inset plots show the convergence rate for the correlation coefficient to ity asymptotic value $c$, always exponential for Tanh and power law for ReLU in the chaotic phase. Solid lines are computed using the IGB approach, while shaded areas represent the 90 $\%$ central confidence interval computed using the MF approach.
@@ 4 Best trainability conditions
Gradients at initialization have extensively been analyzed in the MF literature, so the reader is referred to, for example, [schoenholz2016deep] or [hayou2019impact] for an extensive discussion. For our purposes, it is sufficient to note that $\tilde{\chi}_{1}\equiv\partial q_{ab}^{(l+1)}/\partial q_{ab}^{(l)}|_{c=1}$ is a key quantity separating the ordered phase, where gradients vanish, from the chaotic phase, where gradients explode (App. G). When the signal variance is non-divergent, $\tilde{\chi}_{1}=\chi_{1}$ measures the stability of the fixed point $c=1$, where its dynamic counterpart $c_{ab}^{(l)}$ is connected to $\gamma^{(l)}$ through Eq. (6). In both the ordered phase and at the EOC, the state $c=1$ is a stable fixed point. Consequently, in both cases, we observe an asymptotic $\gamma=\infty$, indicating a state of deep prejudice at initialization.
@@
In [pmlr-v235-francazi24a], one of the main open questions concerned the distinction in dynamical behaviour between neutral and prejudiced phases. The present results not only clarify this distinction by linking it to gradient stability properties, but also reveal a finer structure within the prejudiced phase, identifying conditions that govern the persistence of predictive bias. Moreover, these findings lead to the following conclusion (see proof in App. E), which counters the suggestion of [pmlr-v235-francazi24a], that neutral initializations lead to the fastest dynamics.
@@
From a trainability perspective, the optimal initial condition is not one of neutrality, but rather a state of transient deep prejudice.
@@ 5 Detailed phase diagrams
Due to the equivalence between IGB and MF, all MF results remain valid in the IGB framework. Therefore, for a comprehensive analysis of generic single-node activation functions, we refer to the work of [hayou2019impact]. Beyond that, in App. F we extend the MF/IGB theory to multi-node activation functions, broadening the range of applicability of these theories to e.g. max- and average-pool layers.
@@
Bounded activation functions
In this case, the signal variance is also bounded and the value of $\chi_{1}$ fully delineates the ordered and the chaotic phases. As analyzed in [schoenholz2016deep], the chaotic phase of Tanh — characterized by gradient explosion and training instability — induces a shift of the correlation fixed point to $c<1$, which is shown in Fig. 2 (right plot). Remarkably, the EOC exists for every $\sigma^{2}_{b}\in\mathbb{R}^{+}$ with non-trivial shape [hayou2019impact] (right plot of Fig. 3).
Unbounded activation functions
This case has been extensively analyzed by [hayou2019impact]. However, prior work has overlooked that, for ReLU networks, the correlation coefficient $c^{(l)}$ converges to $c=1$ across the entire phase diagram (see Fig. 2 - left plot), revealing a persistent deep prejudice at initialization. In App. H, we derive explicit recursive relations for the IGB metrics in ReLU networks, demonstrating that $\lim_{l\to\infty}c^{(l)}=1$, while $\gamma^{(l)}$ diverges. Nevertheless, the two MF phases remain distinct, as in the bounded activation case: in the ordered phase, gradients vanish; in the chaotic phase, gradients explode. Crucially, these two phases differ in their depth-scaling behaviour: in the ordered phase, the total signal variance converges and $\gamma^{(l)}$ diverges exponentially with depth, whereas in the chaotic phase, the signal variance diverges and $\gamma^{(l)}$ follows a power-law divergence (Lemma H.1). Hence, persistent deep prejudice may arise via two distinct mechanisms. In the first, the total signal variance ($\sigma^{2}_{y^{(l)}}+\sigma^{2}_{\mu^{(l)}}$) remains bounded while $\sigma^{2}_{y^{(l)}}$ tends toward zero; we denote this phase ordered-deep prejudice, owing to its link with vanishing gradients. In the second mechanism, at least one of $\sigma^{2}_{y^{(l)}}$ or $\sigma^{2}_{\mu^{(l)}}$ diverges, causing network outputs to blow up and gradients to explode. We refer to this as chaotic–deep prejudice. Two independent order parameters are required to distinguish between these regimes.
The behaviour of the signal in the chaotic phase of unbounded activation functions has direct consequences for training with cross entropy loss. Indeed, large separation of output distributions makes the softmax to concentrate entirely to a single class (which depends on the weights initialization), therefore causing a strong per-class dependence of gradients vanishing/exploding (see App. G).
@@ 6 Training dynamics
As an example of more complex architecture, we study empirically a vanilla Vision Transformer [dosovitskiy2020image], where we remove all batch- and layer-norm and skip connections, which may destroy the distinction between phases (indeed, it is easy to show that e.g. a residual MLP has only a critical phase of optimal learning [NIPS2017_81c650ca]). Our simplified Vision Transformer posseses only convolutional, linear and attention layers, and it shows the same transition behaviour of the gradients as the MLP (Fig. 24).
Finally, we validate our assumption by fine-tuning on CIFAR100 a large Vision Transformer (without modifications) pre-trained on ImageNet [deng2009imagenet] (Fig. 4 - bottom). By multiplying all linear and convolutional layers’ weights by a factor $\sigma^{2}_{w}$, we trigger the phase behaviour which we observe in vanilla models. The optimal training state, corresponding to the original non-scaled model, exhibits weak IGB (i.e. maximum classification frequency is significantly greater than $1/n_{c}$, where $n_{c}$ is the number of classes). However, slightly reducing all weights by a factor of $\sigma^{2}_{w}=0.5$ triggers a strong IGB phase, while a larger scaling factor ($\sigma^{2}_{w}=1.5$) reduces IGB, yet hindering the training dynamics. In both cases, the dynamical instabilities encountered either due to vanishing or exploding gradients cannot be absorbed in the early phase of learning.
@@ 7 Conclusions and outlook
In this work, we established an equivalence between two apparently different frameworks for analyzing wide networks at initializaton: mean field theory (MF) and initial guessing bias (IGB). We showed that the fundamental quantities in the two approaches can be mapped onto one another. This connection offered us the possibility to reinterpret the order/chaos phase transition in light of classification bias; the ordered phase is characterized by persistent-deep prejudice, while the chaotic phase is either characterized by persistent-deep prejudice, prejudice or neutrality. Furthermore, our categorization of the edge of chaos (EOC) as a state with deep prejudice reveals that the best trainable model necessarily exhibits bias, which, however, rapidly disappears in the learning dynamics.
Our findings have important implications for understanding the role of architectural choices and hyper-parameter choices in shaping the onset behaviour of deep networks. They suggest that even before training begins, design decisions can inject systematic biases that impact signal propagation, gradient stability, and ultimately trainability. Overall, our work provides a new lens to understand how dataset randomness at initialization, coupled with architectural design, shapes the phase diagram of large MLPs at initialization. We hope this connection between MF and IGB inspires further exploration of the subtle interplay between structure, randomness, and learning dynamics in modern neural networks, as well as of the conditions for reabsorption of initialization biases.
Although experiments in realistic settings are consistent with our theory, a limitation of our work is that to be confined to the infinite-width regime, where the dataset size is brought to infinite afterwards. In this naturally over-parametrized regime, neural networks do not perform feature learning, even though they are capable of generalization (benign overfitting) [lee2018deepneuralnetworksgaussian, 10.1145/3446776]. The extension of our results to the proportional regime, where both the the dataset size and the number of parameters are scaled proportionally to infinite, which is capable of feature learning, remains an interesting, yet unexplored avenue of exploration.
Acknowledgments and Disclosure of Funding
This work was supported by the Swiss National Science Foundation, SNSF grants #\# 196902 and #\# 208249.
@@ Appendix A Extensive notation
  • $\mathbf{\xi}(a))\in\mathbb{R}^{N_{0}}$: $a$-th data sample.
@@
  • $q_{ab}^{ab}{(l)}=\delta_{ij}\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(b)\mkern-1.5mu}\mkern 1.5mu$: the covariance of the pre-activation signals at the $l$-th hidden layer nodes, corresponding to inputs $a$ and $b$, computed before applying the activation function.
@@ Appendix B Training experiments
We train all the models employing the Adam optimizer [kingma2017adammethodstochasticoptimization] without any regularization. The training and test batch sizes are both set to $100$. The learning rate is chosen according to the model and dataset, as reported in Tab. 2. In the next subsections, we report the global accuracy, the accuracy of the most favoured and the most unfavoured classes during training, and the maximum classification frequency, all computed at each training step. In particular, we observe the phase behaviour claimed in Prop. 4.1: phases that are the fastest to train correspond to a maximally biased state, with deep prejudice.
@@ B.2 Residual MLP
We employ a residual MLP studied in the MF regime by [NIPS2017_81c650ca], where the residual branch is rescaled by the total network’s length in order to avoid trivial signal explosion. In particular, [NIPS2017_81c650ca] showed that residual MLP are always critical, i.e. there is no order/chaos phase transition; the correlation coefficient always converges to one at initialization and gradients are always stable (see Fig. 23). In this work, we corroborated this result by training Residual Tanh and ReLU MLPs on binarized Fashion MNIST (Fig. 9 and 10, respectively). Since these models are always critical, the experiments show deep prejudice at beginning of learning, which is absorbed at the same time in all the models, producing high classification accuracies.

### v4

@@ 3 Connecting classification bias to the ordered phase
We report the proof of this theorem in App. E. The key result of Thm. 3.1 establishes a correspondence between the MF formulation and IGB, enabling signal propagation in wide networks to be described interchangeably using either framework. Intriguingly, this result makes no further assumption than the MF regime, and it remains valid for every architecture defined by Eq. (1). In MF theory, $q_{aa}^{(l)}$ and $q_{ab}^{(l)}$ are generally functions of the dataset $\mathcal{D}$, and therefore become random variables upon imposing a distribution over $\mathcal{D}$, as done in the IGB approach. Remarkably, these two quantities concentrate around their mean in the infinite-width and then -data limit, allowing their treatment as deterministic variables (equivalently through $\sigma^{2}_{y^{(l)}}$ and $\sigma^{2}_{\mu^{(l)}}$). On the one hand, this formal connection between MF and IGB frameworks enables a unified view, where predictive behaviour at initialization and trainability conditions are jointly entangled, enriching the classical MF picture, as will be discussed in Secs. 4 and 5. On the other, it extends the IGB framework from pmlr-v235-francazi24a to settings not previously analyzed (e.g., identifying prejudice and neutrality phases for Tanh activations directly from the MF phase diagram (schoenholz2016deep)).
In Fig. 2, we test the validity of Thm. 3.1 by plotting the correlation coefficient in function of the depth for ReLU and Tanh MLPs with $\sigma^{2}_{b}=0.1$ and $\sigma^{2}_{w}$ uniformly varying from the ordered to the chaotic phase analyzed in MF theory. We show similar results on realistic architectures and datasets in App. E.1. In all cases we observe a good agreement between the curves obtained with the IGB approach (solid lines - computed via Eq. (6)) and the 90 $\%$ central confidence interval computed using the MF approach (shaded areas - computed via Eq. (2)). The distribution of MF is very narrow around the IGB-computed values, corroborating the treatment of the signal variance and covariance as deterministic variables. As network’s width increases, the MF distributions of $q_{aa}^{(l)}$ and $q_{ab}^{(l)}$ become progressively more concentrated (see App. E). For ReLU, we observe that the correlation coefficient always converges to one, but the convergence rate is exponential in the ordered phase ($\sigma^{2}_{w}<2$) and follows a power-law in the chaotic phase ($\sigma^{2}_{w}>2$). For Tanh, the correlation coefficient converges to one in the ordered phase and to a lower value in the chaotic phase; in this case, we always observe an exponential convergence behaviour.
@@ 4 Best trainability conditions
Gradients at initialization have extensively been analyzed in the MF literature, so the reader is referred to, for example, schoenholz2016deep or hayou2019impact for an extensive discussion. For our purposes, it is sufficient to note that $\tilde{\chi}_{1}\equiv\partial q_{ab}^{(l+1)}/\partial q_{ab}^{(l)}|_{c=1}$ is a key quantity separating the ordered phase, where gradients vanish, from the chaotic phase, where gradients explode. We show examples of this transition from gradient vanishing to exploding in Fig. 4–left and in App. G). When the signal variance is non-divergent, $\tilde{\chi}_{1}=\chi_{1}$ measures the stability of the fixed point $c=1$, where its dynamic counterpart $c_{ab}^{(l)}$ is connected to $\gamma^{(l)}$ through Eq. (6). In both the ordered phase and at the EOC, the state $c=1$ is a stable fixed point. Consequently, in both cases, we observe an asymptotic $\gamma=\infty$, indicating a state of deep prejudice at initialization.
@@
In pmlr-v235-francazi24a, one of the main open questions concerned the distinction in dynamical behaviour between neutral and prejudiced phases. The present results not only clarify this distinction by linking it to gradient stability properties, but also reveal a finer structure within the prejudiced phase, identifying conditions that govern the persistence of predictive bias. Moreover, these findings lead to the following conclusion (see proof in App. E), which counters the suggestion of pmlr-v235-francazi24a, that neutral initializations lead to the fastest dynamics.
@@
From a trainability perspective, the optimal initial condition (EOC) is not one of neutrality, but rather a state of transient deep prejudice.
@@
How to access the EOC in practice
In real scenarios, when the EOC is not known analytically, one can compute the initialization gradients for different values of $\sigma^{2}_{w}$ and $\sigma^{2}_{b}$ and choose the combination that shows stable gradients across the depth. In App. G, we report initialization gradients for an array of architectures and datasets, showing how the EOC is reached. We also compare the propagation of signals in pretrained vs untrained networks, showing that also in pretrained models one has a similar phenomenology (App. G.4.1).
@@ 5 Detailed phase diagrams
Due to the equivalence between IGB and MF, all MF results remain valid in the IGB framework. Therefore, for a comprehensive analysis of generic single-node activation functions, we refer to the work of hayou2019impact. Beyond that, in App. F we extend the MF/IGB theory to multi-node activation functions, broadening the range of applicability of these theories to e.g. max- and average-pool layers.
@@
Bounded activation functions In this case, the signal variance is also bounded and the value of $\chi_{1}$ fully delineates the ordered and the chaotic phases. As analyzed in schoenholz2016deep, the chaotic phase of Tanh — characterized by gradient explosion and training instability — induces a shift of the correlation fixed point to $c<1$, which is shown in Fig. 2 (right plot). Remarkably, the EOC exists for every $\sigma^{2}_{b}\in\mathbb{R}^{+}$ with non-trivial shape (hayou2019impact) (right plot of Fig. 3).
Unbounded activation functions This case has been extensively analyzed by hayou2019impact. However, prior work has overlooked that, for ReLU networks, the correlation coefficient $c^{(l)}$ converges to $c=1$ across the entire phase diagram (see Fig. 2 - left plot), revealing a persistent deep prejudice at initialization. In App. H, we derive explicit recursive relations for the IGB metrics in ReLU networks, demonstrating that $\lim_{l\to\infty}c^{(l)}=1$, while $\gamma^{(l)}$ diverges. Nevertheless, the two MF phases remain distinct, as in the bounded activation case: in the ordered phase, gradients vanish; in the chaotic phase, gradients explode. Crucially, these two phases differ in their depth-scaling behaviour: in the ordered phase, the total signal variance converges and $\gamma^{(l)}$ diverges exponentially with depth, whereas in the chaotic phase, the signal variance diverges and $\gamma^{(l)}$ follows a power-law divergence (Lemma H.1). Hence, persistent deep prejudice may arise via two distinct mechanisms. In the first, the total signal variance ($\sigma^{2}_{y^{(l)}}+\sigma^{2}_{\mu^{(l)}}$) remains bounded while $\sigma^{2}_{y^{(l)}}$ tends toward zero; we denote this phase ordered-deep prejudice, owing to its link with vanishing gradients. In the second mechanism, at least one of $\sigma^{2}_{y^{(l)}}$ or $\sigma^{2}_{\mu^{(l)}}$ diverges, causing network outputs to blow up and gradients to explode. We refer to this as chaotic–deep prejudice. Two independent order parameters are required to distinguish between these regimes.
Per-class gradient exploding. The behaviour of the signal in the chaotic phase of unbounded activation functions has direct consequences for training with cross entropy loss. Indeed, large separation of output distributions makes the softmax to concentrate entirely to a single class (which depends on the weights initialization), therefore causing a strong per-class dependence of gradients vanishing/exploding (see Fig. 4 and App. G).
Figure 4: Layerwise gradients $\tilde{q}^{(l)}_{ab}$ as a function of the network layer, of a ReLU MLP in the MF regime computed on a batch of CIFAR10. Left: we show the total gradients. Center: we show the most favored class (the one towards which there is most prejudice). Right: the least favored class. In the chaotic phase, the gradients of the favored class are numerically zero. In App. G we explain the notation $\tilde{q}^{(l)}_{ab}$ for the gradients, and show similar figures on different datasets and architectures.
@@ 6 Training dynamics
As an example of more complex architecture, we study empirically a vanilla Vision Transformer (dosovitskiy2020image), where we remove all batch- and layer-norm and skip connections, which may destroy the distinction between phases (indeed, it is easy to show that e.g. a residual MLP has only a critical phase of optimal learning (NIPS2017_81c650ca)). Our simplified Vision Transformer posseses only convolutional, linear and attention layers, and it shows the same transition behaviour of the gradients as the MLP (Fig. 27).
Finally, we validate our assumption by fine-tuning on CIFAR100 a large Vision Transformer (without modifications) pre-trained on ImageNet (deng2009imagenet) (Fig. 5 - bottom). By multiplying all linear and convolutional layers’ weights by a factor $\sigma^{2}_{w}$, we trigger the phase behaviour which we observe in vanilla models. The optimal training state, corresponding to the original non-scaled model, exhibits weak IGB (i.e. maximum classification frequency is significantly greater than $1/n_{c}$, where $n_{c}$ is the number of classes). However, slightly reducing all weights by a factor of $\sigma^{2}_{w}=0.5$ triggers a strong IGB phase, while a larger scaling factor ($\sigma^{2}_{w}=1.5$) reduces IGB, yet hindering the training dynamics. In both cases, the dynamical instabilities encountered either due to vanishing or exploding gradients cannot be absorbed in the early phase of learning.
@@ 7 Conclusions and outlook
We established an equivalence between two independent frameworks for the analysis of wide networks at initializaton: mean field theory (MF) and initial guessing bias (IGB). We showed that the fundamental quantities in the two approaches can be mapped onto one another. This connection offered us the possibility to reinterpret the order/chaos phase transition in light of classification bias; the ordered phase is characterized by persistent-deep prejudice, while the chaotic phase is either characterized by persistent-deep prejudice, prejudice or neutrality. Furthermore, our categorization of the edge of chaos (EOC) as a state with deep prejudice reveals that the best trainable model necessarily exhibits bias, which, however, rapidly disappears in the learning dynamics.
Our findings have important implications for understanding the role of architectural choices and hyper-parameter choices in shaping the onset behaviour of deep networks. They suggest that even before training begins, design decisions can inject systematic biases that impact signal propagation, gradient stability, and ultimately trainability.
Overall, our work provides a new lens to understand how dataset randomness at initialization, coupled with architectural design, shapes the phase diagram of large MLPs at initialization. We hope this connection between MF and IGB inspires further exploration of the subtle interplay between structure, randomness, and learning dynamics in modern neural networks, as well as of the conditions for reabsorption of initialization biases.
Practical Takewaways Finally, we highlight two practical takeaways of our work. First: neutral initializations are inefficient, so typical initializations are prejudiced. This means that, in order to be meaningful, hyperparameter tuning runs need to be long enough to reabsorb IGB. As our runs show, the EOC is not only optimal from the point of view of stability of the gradients, but also in terms of how fast the biases are reabsorbed. A consequence is that, if one tunes the weights’ variances in order to be at the edge of chaos (as explained in Sec. 4), shorter HP tuning runs are needed. Second: since initializations are biased, the gradients of the favored classes will typically be smaller than those of the unfavored. This is especially true with gradient exploding in the chaotic deep prejudice phase, where one class has exploding, while the other(s) have zero gradients (Fig. 4). Such disparities in the per-class gradients can have drastical impacts on the learning speed and quality, as demonstrated in francazi2023theoretical.
Reproducibility statement
We provide the full code used to replicate all the simulations and training experiments at the link: https://github.com/abassi98/igb_and_trainability.git. We declare the usage of Large Language Models for finding related work.
Acknowledgements
This work was supported by the Swiss National Science Foundation, SNSF grants # 208249 and # 196902.
@@ Appendix A Extensive notation
  • $\mathbf{\xi}(a)\in\mathbb{R}^{N_{0}}$: $a$-th data sample.
@@
  • $q_{ab}^{(l)}=\delta_{ij}\mkern 1.5mu\overline{\mkern-1.5muy_{i}^{(l)}(a)y_{j}^{(l)}(b)\mkern-1.5mu}\mkern 1.5mu$: the covariance of the pre-activation signals at the $l$-th hidden layer nodes, corresponding to inputs $a$ and $b$, computed before applying the activation function.
@@ Appendix B Training experiments
We train all the models employing the Adam optimizer (kingma2017adammethodstochasticoptimization) without any regularization. The training and test batch sizes are both set to $100$. The learning rate is chosen according to the model and dataset, as reported in Tab. 2. In the next subsections, we report the global accuracy, the accuracy of the most favoured and the most unfavoured classes during training, and the maximum classification frequency, all computed at each training step. In particular, we observe the phase behaviour claimed in Prop. 4.1: phases that are the fastest to train correspond to a maximally biased state, with deep prejudice.
@@ B.2 Residual MLP
We employ a residual MLP studied in the MF regime by NIPS2017_81c650ca, where the residual branch is rescaled by the total network’s length in order to avoid trivial signal explosion. In particular, NIPS2017_81c650ca showed that residual MLP are always critical, i.e. there is no order/chaos phase transition; the correlation coefficient always converges to one at initialization and gradients are always stable (see Fig. 26). In this work, we corroborated this result by training Residual Tanh and ReLU MLPs on binarized Fashion MNIST (Fig. 10 and 11, respectively). Since these models are always critical, the experiments show deep prejudice at beginning of learning, which is absorbed at the same time in all the models, producing high classification accuracies.